Testing

John Doe

Abstract

This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test. This is a test.

* Installation Changes in Emacs 21.1

See the INSTALL file for information on installing extra libraries and fonts to take advantage of the new graphical features and extra charsets in this release.

** Support for GNU/Linux on IA64 machines has been added.

** Support for LynxOS has been added.

** There are new configure options associated with the support for images and toolkit scrollbars. Use the –help option in ‘configure’ to list them.

** You can build a 64-bit Emacs for SPARC/Solaris systems which support 64-bit executables and also on Irix 6.5. This increases the maximum buffer size. See etc/MACHINES for instructions. Changes to build on other 64-bit systems should be straightforward modulo any necessary changes to unexec.

** There is a new configure option ‘–disable-largefile’ to omit Unix-98-style support for large files if that is available.

** There is a new configure option ‘–without-xim’ that instructs Emacs to not use X Input Methods (XIM), if these are available.

** ‘movemail’ defaults to supporting POP. You can turn this off using the –without-pop configure option, should that be necessary.

** This version can be built for the Macintosh, but does not implement all of the new display features described below. The port currently lacks unexec, asynchronous processes, and networking support. See the ”Emacs and the Mac OS” appendix in the Emacs manual, for the description of aspects specific to the Mac.

** Note that the MS-Windows port does not yet implement various of the new display features described below.